OpenStudy (anonymous):

Determine each function as even, odd, or neither. Prove it! (#21) f(x)= sinx (22.) f(x)=cos2x (23.) f(x)= sec x (24.)f(x)=x^3+1 (25.)f(x)=3sq. root 2x (26.)f(x)=x^3+2x

OpenStudy (anonymous):

sinx is even, because sin(pi/2)= 1 and sin(-pi/2)=-1

OpenStudy (anonymous):

sinx is odd

OpenStudy (anonymous):

i think you might have the definitions mixed up. a function is even if: \[f(-x) = f(x)\] and odd if: \[f(-x) = -f(x)\] sin(x) is odd because: \[s(-x) = -\sin(x)\]

OpenStudy (anonymous):

cos 2x should be even

OpenStudy (anonymous):

secx even

OpenStudy (anonymous):

3sqrt(2x) can't be negative so maybe not even and not odd

OpenStudy (anonymous):

rest should be odd

OpenStudy (anonymous):

x^3+1 is neither. x^3 by itself is odd, the +1 throws things off.

OpenStudy (anonymous):

Yes thanks Joe................... I forgot the 1 effect lol